OpenStudy (anonymous):

What is the formula to convert meter per secound into meter per hour(m/s into m/hr)

OpenStudy (anonymous):

\[1hr = 60 \times 60 seconds\] 1 hr = 3600 seconds 1 second= (1/3600 ) hours m/s = m/ [ (1/3600) hrs ] m/s = 3600 m/hr